OpenStudy (lovelyharmonics):

Write the expression in standard form. five divided by quantity three minus fifteen i.

OpenStudy (lovelyharmonics):

(5/78)-(25/78)i (5/78)+(25/78)i -(5/78)+(25/78)i -(5/78)-(25/78)i

OpenStudy (lovelyharmonics):

it has to be one of those .-. @surjithayer

OpenStudy (anonymous):

\[\frac{ 5 }{ 3-15\iota }\times \frac{ 3+15\iota }{ 3+15\iota }\] simplify it. \[use~ (a+b \iota )(a-b \iota )=a^2-\iota^2b^2=a^2+b^2\]

OpenStudy (lovelyharmonics):

where do you find the beginning of that equation?

OpenStudy (anonymous):

previously i have not seen i after fifteen

OpenStudy (lovelyharmonics):

okay, but i don't understand where to put everything in that equation

OpenStudy (anonymous):

i have rationalise the denominator.

OpenStudy (anonymous):

multiply the numerator,denominator i have shown above to solve it.

OpenStudy (lovelyharmonics):

sorry my computer shut down .-. okay so to do that would you foil the denominators and numerators with eachother?

OpenStudy (anonymous):

\[=\frac{ 5\left( 3+15\iota \right) }{ \left( 3 \right)^2-\left( 15\iota \right)^2 }\] \[=\frac{ 15+75\iota }{9-225i^2 }\] \[\iota^2=-1\] now you can solve.

OpenStudy (lovelyharmonics):

(15+75i)/134 15/134 +75i/134 but how do you simplify 134 into 78? @surjithayer

OpenStudy (anonymous):

it is 234 not 134 cancel the numerator and denominator by 3

OpenStudy (anonymous):

9+225=234

OpenStudy (lovelyharmonics):

3/78 + 25i/78 so they would both be positive?

OpenStudy (anonymous):

5/78 not 3/78

OpenStudy (lovelyharmonics):

oh okay thank you ^.^

OpenStudy (anonymous):

yw